# Datalog-Expressibility for Monadic and Guarded Second-Order Logic

### v2

@@ Abstract
We characterise the sentences in Monadic Second-order Logic (MSO) that are over finite structures equivalent to a Datalog program, in terms of an existential pebble game. We also show that for every class ${\mathcal{C}}$ of finite structures that can be expressed in MSO and is closed under homomorphisms, and for all $\ell,k\in{\mathbb{N}}$, there exists a canonical Datalog program $\Pi$ of width $(\ell,k)$ in the sense of Feder and Vardi. The same characterisations also hold for Guarded Second-order Logic (GSO), which properly extends MSO. To prove our results, we show that every class ${\mathcal{C}}$ in GSO whose complement is closed under homomorphisms is a finite union of constraint satisfaction problems (CSPs) of $\omega$-categorical structures. The intersection of MSO and Datalog is known to contain the class of nested monadically defined queries (Nemodeq); likewise, we show that the intersection of GSO and Datalog contains all problems that can be expressed by the more expressive language of nested guarded queries (GQ+). Yet, by exploiting our results, we can show that neither of the two query languages can serve as a characterization, as we exhibit a query in the intersection of MSO and Datalog that is not expressible in GQ+.
@@ 1 Introduction
Another fundamental formalism in theoretical computer science, which is particularly heavily studied in database theory and logic-based knowledge representation, is Datalog (see, e.g., [26, 38]). Every fixed Datalog program can be evaluated on given finite structures in polynomial time. If a linear order is preseent and if the available relations are closed under complement, Datalog can even express all polytime-computable queries [1]. That is, like MSO, Datalog strikes a good balance between expressivity and good mathematical and computational properties.
Neither of the two formalisms subsumes the other in terms of what can be expressed. Yet, in various scenarios, we are interested in simultaneously having the good computational properties of expressibility in Datalog and having the good computational properties of expressibility in MSO or GSO. A wide variety of popular query formalisms (among them (unions of) conjunctive queries, (2-way conjunctive) regular path queries, monadic Datalog, the recently introduced almost monadic Datalog  [43], guarded Datalog, monadically defined queries, or nested monadically defined queries) are known to be both in Datalog and GSO [42]. Query languages expressible both in Datalog and MSO have been shown to warrant decidable query entailment over logical theories exhibiting universal models of finite treewidth and cliquewidth [32, 31]. Also, all these formalisms have favourable properties when it comes to static analysis, most notably decidable query containment [42]. Note that on the contrary, query containment in unrestricted Datalog is undecidable, as is query containment in unrestricted MSO / GSO. The same holds for query entailment under logical theories, even if the latter are extremely simple. So it is really the interplay of the restrictions imposed by both formalisms that is required to ensure decidability of central tasks in databases and knowledge representation. This makes the semantic intersection of Datalog and MSO / GSO so interesting and worthwhile investigating.
In this article, we investigate two questions that turn out to be closely related:
@@
The constraint satisfaction problem for the structure $({\mathbb{Q}};<)$, which is the class of all finite acyclic digraphs $(V;E)$, cannot be expressed in MMSNP [8]. It can, however, be expressed in MSO by the sentence
@@
To see this, note that if $(V;E)$ is a digraph such that there exists a non-empty $X\subseteq V$ such that for every $x\in X$ there is $y\in Z$ with $E(x,y)$, then $X$ contains a directed cycle, and hence has no homomorphism to $({\mathbb{Q}};<)$. Conversely, if $(V;E)$ is a finite digraph and contains no directed cycle, then every non-empty subset of $V$ must contain a sink, i.e., a vertex $x$ with no outgoing $E$-edges, and hence satisfies the given sentence.
The class of CSPs of arbitrary infinite structures $\mathfrak{B}$ is quite large; it is easy to see that, given any finite relational signature $\tau$, a class $\mathcal{D}$ of finite $\tau$-structures is a CSP of some countably infinite structure if and only if
@@
The second condition, which is sometimes also referred to as closure under inverse homomorphisms, can be equivalently rephrased by requiring that the complement [^1] Whenever in this paper we refer to the complement of some class $\mathcal{D}$ of finite $\tau$-structures, we mean the class of all finite $\tau$-structures not contained in $\mathcal{D}$ – we will make $\tau$ explicit or it will be clear from the context. of ${\mathcal{D}}$ is closed under homomorphisms, where a class $\mathcal{C}$ is closed under homomorphisms if for any structure $\mathfrak{A}\in\cal C$ that maps homomorphically to some $\mathfrak{C}$ we have $\mathfrak{C}\in\cal C$. Examples of classes of structures that are closed under homomorphisms naturally arise from Datalog. We say that a class $\mathcal{C}$ of finite $\tau$-structures is definable in Datalog [^2] Warning: Feder and Vardi [29] say that a CSP is in Datalog if its complement in the class of all finite $\tau$-structures is in Datalog. if there exists a Datalog program $\Pi$ with a distinguished nullary predicate ${\sf goal}$  such that $\Pi$ derives ${\sf goal}$  on a finite $\tau$-structure if and only if the structure is in $\mathcal{C}$; in this case, we may denote $\mathcal{C}$ by $\llbracket\Pi\rrbracket$. Every class of $\tau$-structures in Datalog is closed under homomorphisms, but the converse does not hold: there are classes of finite structures in Datalog not corresponding to the complement of any CSP.
@@
does not derive ${\sf goal}$  on that structure.
A noteworthy subclass of all CSPs are the CSPs of structures $\mathfrak{B}$ that are countably infinite and $\omega$-categorical. A structure $\mathfrak{B}$ is $\omega$-categorical if all countable models of $\mathfrak{B}$’s first-order theory are isomorphic to $\mathfrak{B}$. A well-known example of an $\omega$-categorical structure is $({\mathbb{Q}};<)$, a result going back to Cantor [20]. Constraint satisfaction problems of $\omega$-categorical structures can be evaluated in polynomial time on any class of structures whose treewidth is bounded by some constant $k\in{\mathbb{N}}$, by a result of Bodirsky and Dalmau [12], using Datalog programs.
Two important parameters of a Datalog program $\Pi$ are the maximal arity $\ell$ of its auxiliary predicates (IDBs), and the maximal number $k$ of variables per rule in $\Pi$ – we then say that $\Pi$ has width $(\ell,k)$, following the terminology of Feder and Vardi [29]. The special case of $\ell=1$ is typically referred to as monadic Datalog. The parameters $\ell$ and $k$ are important in theory, but they also bear some practical relevance: When evaluating $\Pi$ on a given structure $\mathfrak{A}$ with domain $A$, the memory needed is bounded by $O(|A|^{\ell})$ and the number of computation steps by $O(|A|^{k})$. The polynomial-time algorithm for instances of an $\omega$-categorical CSP of treewidth at most $k$ presented by Bodirsky and Dalmau is in fact a Datalog program of width $(k-1,k)$. A Datalog program $\Pi$ is called sound for a class of $\tau$-structures ${\mathcal{C}}$ if $\llbracket\Pi\rrbracket\subseteq{\mathcal{C}}$. Bodirsky and Dalmau showed that if ${\mathcal{C}}$ is the complement of the CSP of an $\omega$-categorical $\tau$-structure $\mathfrak{B}$ then there exists, for any $\ell,k\in{\mathbb{N}}$, a Datalog program $\Pi$ of width $(\ell,k)$ such that
@@
Interestingly and conveniently, there is a game-theoretic characterisation capturing whether the canonical Datalog program of width $(\ell,k)$ for ${\mathcal{C}}$ derives ${\sf goal}$ on a given $\tau$-structure $\mathfrak{A}$ [12]. This characterisation is based on the existential pebble game from finite model theory, which is played on a pair $(\mathfrak{A},\mathfrak{B})$ of structures. In more detail, the existential $(\ell,k)$-pebble game is played by two players, called Spoiler and Duplicator (see, e.g., [25, 29, 35]). Spoiler starts by placing $k$ pebbles on elements $a_{1},\dots,a_{k}$ of $\mathfrak{A}$, and Duplicator responds by placing $k$ pebbles $b_{1},\dots,b_{k}$ on $\mathfrak{B}$. If the map that sends $a_{1},\dots,a_{k}$ to $b_{1},\dots,b_{k}$ is not a partial homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$, then the game is over and Spoiler wins. Otherwise, Spoiler removes all but at most $\ell$ pebbles from $\mathfrak{A}$, and Duplicator has to respond by removing the corresponding pebbles from $\mathfrak{B}$. Then Spoiler can again place all his pebbles on $\mathfrak{A}$, and Duplicator must again respond by placing her pebbles on $\mathfrak{B}$. If the game continues forever, then Duplicator wins. If $\mathfrak{B}$ is a finite, or more generally a countable $\omega$-categorical structure then Spoiler has a winning strategy for the existential $(\ell,k)$-pebble game on $(\mathfrak{A},\mathfrak{B})$ if and only if the canonical Datalog program for $\operatorname{CSP}(\mathfrak{B})$ derives ${\sf goal}$ on $\mathfrak{A}$ (Theorem 4.2). This connection played an essential role in proving Datalog inexpressibility results, for example for the class of finite-domain CSPs [3] (leading to a complete classification of those finite structures $\mathfrak{B}$ such that the complement of $\operatorname{CSP}(\mathfrak{B})$ can be expressed in Datalog [4]).
@@
In this article, we present a characterisation of those GSO sentences $\Phi$ that are over finite structures equivalent to a Datalog program. Our characterisation involves a variant of the existential pebble game from finite model theory, which we call the $(\ell,k)$-game. This game is defined for a homomorphism-closed class ${\mathcal{C}}$ of finite $\tau$-structures, and it is played by the two players Spoiler and Duplicator on a finite $\tau$-structure $\mathfrak{A}$ as follows.
@@
In Section 4 we show that a GSO sentence $\Phi$ is over finite structures equivalent to a Datalog program of width $(\ell,k)$ if and only if
@@
We also show that for every GSO sentence $\Phi$ whose class of finite models ${\mathcal{C}}$ is closed under homomorphisms and for all $\ell,k\in{\mathbb{N}}$ there exists a canonical Datalog program $\Pi$ of width $(\ell,k)$ for $\mathcal{C}$ (Theorem 4.12). To prove these results, we first show that every class of finite structures in GSO whose complement is closed under homomorphisms is a finite union of CSPs that can also be expressed in GSO (Lemma 3.17; an analogous statement holds for MSO). Moreover, every CSP in GSO is the CSP of a countable $\omega$-categorical structure (Corollary 3.8); this allows us to use results from [12] to make the link to existential pebble games. We highlight that this result elegantly generalises known results:
Note that our results imply that every class of finite structures that can be expressed both in GSO and in Datalog is an intersection of finitely many complemented CSPs of $\omega$-categorical structures. In contrast, it is not generally true that a Datalog program describes a finite intersection of complements of CSPs (we present a counterexample in Example 3.20). We also present an example of a CSP which is expressible in MSO and coNP-complete, and hence not the CSP of a reduct of a finitely bounded homogeneous structure, unless NP=coNP (Proposition 5.1).
Our final results concern the most expressive syntactically defined formalism known from the literature that is contained in both Datalog and MSO, namely nested monadically defined queries [42], and the most expressive syntactically defined formalism known from the literature that is contained in both Datalog and GSO, namely nested guarded queries [18].
We prove that there are problems in the intersection of Datalog and GSO that cannot be expressed as a nested guarded query. To prove this result, we introduce a modified version of the existential pebble game, which we call the nested guarded pebble game, and which captures precisely the expressiveness of nested guarded queries (Theorem 6.26). We also present an example of a problem which lies in the intersection of Datalog and MSO that cannot be expressed by nested monadically defined queries (and not even by nested guarded queries; Corollary 6.51).
@@ 2 Preliminaries
It is easy to see that every existential positive $\tau$-formula is a disjunction of primitive positive $\tau$-formulas (and hence referred to as a union of conjunctive queries in database theory). If $\Phi$ is a primitive positive $\tau$-formula without equality for a relational signature $\tau$, then the canonical database is the $\tau$-structure $\mathfrak{A}$ whose domain consists of all the variables of $\Phi$, and where $R^{\mathfrak{A}}$, for $R\in\tau$ of arity $k$, consists of all tuples $(x_{1},\dots,x_{k})$ such that $\Phi$ contains the conjunct $R(x_{1},\dots,x_{k})$.
Formulas without free variables are called sentences; in database theory, formulas are often called queries and sentences are often called Boolean queries. If $\Phi$ is a sentence, we write $\llbracket\Phi\rrbracket$ for the class of all finite models of $\Phi$.
A reduct of a relational structure $\mathfrak{A}$ is a structure $\mathfrak{A}^{\prime}$ obtained from $\mathfrak{A}$ by dropping some of the relations, and $\mathfrak{A}$ is called an expansion of $\mathfrak{A}^{\prime}$.

### v3

@@ Abstract
We characterise the sentences in Monadic Second-order Logic (MSO) that are over finite structures equivalent to a Datalog program, in terms of an existential pebble game. We also show that for every class $\mathcal{C}$} of finite structures that can be expressed in MSO and is closed under homomorphisms, and for all $\ell,k\in{\mathbb{N}}$, there exists a canonical Datalog program $\Pi$ of width $(\ell,k)$ in the sense of Feder and Vardi. The same characterisations also hold for Guarded Second-order Logic (GSO), which properly extends MSO. To prove our results, we show that every class $\mathcal{C}$} in GSO whose complement is closed under homomorphisms is a finite union of constraint satisfaction problems (CSPs) of $\omega$-categorical structures. The intersection of MSO and Datalog is known to contain the class of nested monadically defined queries (Nemodeq); likewise, we show that the intersection of GSO and Datalog contains all problems that can be expressed by the more expressive language of nested guarded queries (GQ+). Yet, by exploiting our results, we can show that neither of the two query languages can serve as a characterization, as we exhibit a CSP whose complement corresponds to a query in the intersection of MSO and Datalog that is not expressible in GQ+.
@@ 1 Introduction
Another fundamental formalism in theoretical computer science, which is particularly heavily studied in database theory and logic-based knowledge representation, is Datalog (see, e.g., [29, 40]). Every fixed Datalog program can be evaluated on given finite structures in polynomial time. If a linear order is present and if the available relations are closed under complement, Datalog can even express all polytime-computable queries [1]. That is, like MSO and GSO, Datalog strikes a good balance between expressivity and good mathematical and computational properties.
Neither of the two formalisms subsumes the other in terms of what can be expressed. Yet, in various scenarios, we are interested in simultaneously having the good computational properties of expressibility in Datalog and having the good computational properties of expressibility in MSO or GSO. A wide variety of popular query formalisms (among them (unions of) conjunctive queries, (2-way conjunctive) regular path queries, monadic Datalog, the recently introduced almost monadic Datalog [44], guarded Datalog, monadically defined queries, or nested monadically defined queries) are known to be both in Datalog and GSO [43]. Query languages expressible both in Datalog and MSO have been shown to warrant decidable query entailment over logical theories exhibiting universal models of finite treewidth and cliquewidth [35, 34]. Also, all these formalisms have favourable properties when it comes to static analysis, most notably decidable query containment [43, 21]. Note that on the contrary, query containment in unrestricted Datalog is undecidable, as is query containment in unrestricted MSO / GSO [^1] This follows from Trakhtenbrot’s theorem which states that satisfiability of first-order logic in the finite is undecidable.. The same holds for query entailment under logical theories, even if the latter are extremely simple. So it is really the interplay of the restrictions imposed by both formalisms that is required to ensure decidability of central tasks in databases and knowledge representation. This makes the semantic intersection of Datalog and MSO / GSO so interesting and worthwhile investigating.
In this article, we investigate two questions, which will turn out to be closely related:
@@
The constraint satisfaction problem for the structure $({\mathbb{Q}};<)$, which is the class of all finite acyclic digraphs $(V;E)$, cannot be expressed in MMSNP [10]. It can, however, be expressed in MSO over the class of all finite digraphs by the sentence
@@
To see this, note that if $(V;E)$ is a finite digraph such that there exists a non-empty $X\subseteq V$ such that for every $x\in X$ there is a $y\in Z$ with $E(x,y)$, then $X$ contains a directed cycle, and hence has no homomorphism to $({\mathbb{Q}};<)$. Conversely, if $(V;E)$ is a finite digraph and contains no directed cycle, then every non-empty subset of $V$ must contain a sink, i.e., a vertex $x$ with no outgoing $E$-edges, and hence satisfies the given sentence. $\triangle$
The class of CSPs of arbitrary infinite structures $\mathfrak{B}$ is quite large; the following is easy to see (see, e.g., [10, Lemma 1.1.8]).
Remark 1.2.
For every finite relational signature $\tau$, a class $\mathcal{D}$ of finite $\tau$-structures is the CSP of some countably infinite structure if and only if
@@
A class of finite relational structures that satisfies these two conditions is simply called a CSP.
The second condition in Remark 1.2, which is sometimes also referred to as closure under inverse homomorphisms, can be equivalently rephrased by requiring that the complement [^2] Whenever in this paper we refer to the complement of some class $\mathcal{D}$ of finite $\tau$-structures, we mean the class of all finite $\tau$-structures not contained in $\mathcal{D}$ – we will make $\tau$ explicit or it will be clear from the context. of $\mathcal{D}$} is closed under homomorphisms, where a class $\mathcal{C}$ is closed under homomorphisms if for any structure $\mathfrak{A}\in\cal C$ that maps homomorphically to some $\mathfrak{C}$ we have $\mathfrak{C}\in\cal C$. Examples of classes of structures that are closed under homomorphisms naturally arise from Datalog. We say that a class $\mathcal{C}$ of finite $\tau$-structures is definable in Datalog (or shorter: is in Datalog) [^3] Warning: Feder and Vardi [32] say that a CSP is in Datalog if its complement in the class of all finite $\tau$-structures is definable in Datalog in our sense. if there exists a Datalog program $\Pi$ with a distinguished nullary predicate $\sf goal$} such that $\Pi$ derives $\sf goal$} on a finite $\tau$-structure if and only if the structure is in $\mathcal{C}$; in this case, we may denote $\mathcal{C}$ by $\llbracket\Pi\rrbracket$. While each class of $\tau$-structures that is definable in Datalog is closed under homomorphisms, not every such class corresponds to the complement of a CSP.
@@
does not derive $\sf goal$} on that structure. $\triangle$
A noteworthy subclass of all CSPs are the CSPs of structures $\mathfrak{B}$ that are countably infinite and $\omega$-categorical. A structure $\mathfrak{B}$ is $\omega$-categorical if all countable models of $\mathfrak{B}$’s first-order theory are isomorphic to $\mathfrak{B}$. A well-known example of an $\omega$-categorical structure is $({\mathbb{Q}};<)$, a result going back to Cantor [23]. We would like to mention that constraint satisfaction problems of $\omega$-categorical structures can be evaluated in polynomial time on any class of structures whose treewidth is bounded by some constant $k\in{\mathbb{N}}$, by a result of Bodirsky and Dalmau [14], using Datalog programs; however, the concept of bounded treewidth is not needed in the present article.
Two important parameters of a Datalog program $\Pi$ are the maximal arity $\ell$ of its auxiliary predicates (IDBs), and the maximal number $k$ of variables per rule in $\Pi$ – we then say that $\Pi$ has width $(\ell,k)$, following the terminology of Feder and Vardi [32]. The special case of $\ell=1$ is typically referred to as monadic Datalog. The parameters $\ell$ and $k$ are important in theory, but they also bear some practical relevance: When evaluating $\Pi$ on a given structure $\mathfrak{A}$ with domain $A$, the memory needed is bounded by $O(|A|^{\ell})$ and the number of computation steps by $O(|A|^{k})$. The polynomial-time algorithm for instances of an $\omega$-categorical CSP of treewidth at most $k$ presented by Bodirsky and Dalmau is in fact a Datalog program of width (k1,k)(k-1,k). A Datalog program $\Pi$ is called sound for a class of $\tau$-structures $\mathcal{C}$} if $\llbracket\Pi\rrbracket\subseteq{\mathcal{C}}$. Bodirsky and Dalmau showed that if $\mathcal{C}$} is the complement of the CSP of an $\omega$-categorical $\tau$-structure $\mathfrak{B}$ then there exists, for any $\ell,k\in{\mathbb{N}}$, a Datalog program $\Pi$ of width $(\ell,k)$ such that
@@
Interestingly and conveniently, there is a known game-theoretic characterisation capturing whether the canonical Datalog program of width $(\ell,k)$ for $\mathcal{C}$} derives $\sf goal$} on a given $\tau$-structure $\mathfrak{A}$ [14]. This characterisation is based on the existential pebble game from finite model theory, which is played on a pair $(\mathfrak{A},\mathfrak{B})$ of structures. In more detail, the existential $(\ell,k)$-pebble game is played by two players, called Spoiler and Duplicator (see, e.g., [28, 32, 38]). Spoiler starts by placing $k$ pebbles on elements $a_{1},\dots,a_{k}$ of $\mathfrak{A}$, and Duplicator responds by placing $k$ pebbles $b_{1},\dots,b_{k}$ on $\mathfrak{B}$. If the map that sends $a_{1},\dots,a_{k}$ to $b_{1},\dots,b_{k}$ is not a partial homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$, then the game is over and Spoiler wins. Otherwise, Spoiler removes all but at most $\ell$ pebbles from $\mathfrak{A}$, and Duplicator has to respond by removing the corresponding pebbles from $\mathfrak{B}$. Then Spoiler can again place all his pebbles on $\mathfrak{A}$, and Duplicator must again respond by placing her pebbles on $\mathfrak{B}$. If the game continues forever, then Duplicator wins. If $\mathfrak{B}$ is a finite, or more generally a countable $\omega$-categorical structure then Spoiler has a winning strategy for the existential $(\ell,k)$-pebble game on $(\mathfrak{A},\mathfrak{B})$ if and only if the canonical Datalog program of width $(\ell,k)$ for $\operatorname{CSP}(\mathfrak{B})$ derives $\sf goal$} on $\mathfrak{A}$ (Theorem 4.2). This connection played an essential role in proving Datalog inexpressibility results, for example for the class of finite-domain CSPs [3] (leading to a complete classification of those finite structures $\mathfrak{B}$ such that the complement of $\operatorname{CSP}(\mathfrak{B})$ can be expressed in Datalog [5]).
@@
In this article, we prove that every class of finite structures in GSO whose complement is closed under homomorphisms is a finite union of CSPs that can also be expressed in GSO (Theorem 3.22; an analogous statement holds for MSO). Moreover, every CSP in GSO is the CSP of a countable $\omega$-categorical structure (Corollary 3.10). We highlight that this result elegantly generalises known results:
It follows that every class of finite structures that can be expressed both in GSO and in Datalog is an intersection of finitely many complemented CSPs of $\omega$-categorical structures. In contrast, it is not generally true that a Datalog program describes a finite intersection of complements of CSPs (we present a counterexample in Example 3.25).
Next, we use the connection between GSO and CSPs to present a characterisation of those GSO sentences $\Phi$ that are over finite structures equivalent to a Datalog program (Section 4). Our characterisation involves a variant of the existential pebble game from finite model theory, which we call the $(\ell,k)$-game. This game is defined for a homomorphism-closed class $\mathcal{C}$} of finite $\tau$-structures, and it is played by the two players Spoiler and Duplicator on a finite $\tau$-structure $\mathfrak{A}$ as follows.
@@
We use results from [14] to show that a GSO sentence $\Phi$ is over finite structures equivalent to a Datalog program of width $(\ell,k)$ if and only if
@@
We also show that for every class of finite models $\mathcal{C}$} which is closed under homomorphisms and expressible in GSO, and for all $\ell,k\in{\mathbb{N}}$, there exists a canonical Datalog program $\Pi$ of width $(\ell,k)$ for $\mathcal{C}$ (Theorem 4.12).
Our next series of results concern the most expressive syntactically defined formalism known from the literature that is contained in both Datalog and MSO, namely nested monadically defined queries [43], as well as the most expressive syntactically defined formalism known from the literature that is contained in both Datalog and GSO, namely nested guarded queries [21].
We prove that there are problems in the intersection of Datalog and GSO that cannot be expressed as a nested guarded query (Section 5). To prove this result, we introduce a modified version of the existential pebble game, which we call the nested guarded pebble game, and which captures precisely the expressiveness of nested guarded queries (Theorem 5.27). We also present an example of a problem which lies in the intersection of Datalog and MSO that cannot be expressed by nested monadically defined queries (and not even by nested guarded queries; Corollary 5.39).
Finally, we provide evidence that the class of CSPs that can be expressed in MSO is not contained in the intensively studied class of CSPs for reducts of finitely bounded homogeneous structures: in Section 6 we present an example of a CSP which is expressible in MSO and coNP-complete, and hence not the CSP of a reduct of a finitely bounded homogeneous structure, unless NP=coNP (Proposition 6.4). As an illustration of the results of Section 4, we also prove that the given MSO sentence is not equivalent to a Datalog program.
@@ 2 Preliminaries
It is easy to see that every existential positive $\tau$-formula is a disjunction of primitive positive $\tau$-formulas (and hence referred to as a union of conjunctive queries in database theory). If $\phi$ is a primitive positive $\tau$-formula without equality for a relational signature $\tau$, then the canonical database is the $\tau$-structure $\mathfrak{A}$ whose domain consists of all the variables of $\phi$, and where $R^{\mathfrak{A}}$, for $R\in\tau$ of arity $k$, consists of all tuples $(x_{1},\dots,x_{k})$ for which $\phi$ contains the conjunct $R(x_{1},\dots,x_{k})$.
Formulas without free variables are called sentences; in database theory, formulas are often called queries and sentences are often called Boolean queries. If $\phi$ is a sentence, we let $\llbracket\phi\rrbracket$ denote the class of all finite models of $\phi$.
A reduct of a relational structure $\mathfrak{A}$ is a structure $\mathfrak{A}^{\prime}$ obtained from $\mathfrak{A}$ by dropping some of the relations. $\mathfrak{A}$ is then also called an expansion of $\mathfrak{A}^{\prime}$.